OpenStudy (anonymous):

Sandra is riding the Ferris wheel, and her height can be modeled by the equation H(t) = 25 cospi over 14t + 31, where H represents the height of the person above the ground in feet at t seconds.

OpenStudy (anonymous):

Part 1: How far above the ground is Sandra before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Sandra begins the ride at the top, how far from the ground is the edge of the Ferris wheel when Sandra's height above the ground reaches a minimum?

OpenStudy (aric200):

Part 1: At t = 0 is when the ride begins. H(0) = 25cos(0) + 31 = 25 + 31 = 56 feet Part 2: When the wheel makes one complete revolution, Sandra will be at her original height when the ride started(56 feet). H(t) = 25*cos((pi/14)t) + 31 = 56 cos((pi/14)t) = 1 (pi/14)t = 2pi => t = 28 seconds Part 3: Minimum occurs when t = 14 seconds. The height at the time will be H(14) = 31 - 25 = 6 feet

OpenStudy (anonymous):

Ya, I found that but I don't know if the first part is correct. Can you explain?

OpenStudy (anonymous):

@Hero @jim_thompson5910 @myininaya

OpenStudy (anonymous):

@wio

jimthompson5910 (jim_thompson5910):

does aric200's answer make sense?

jimthompson5910 (jim_thompson5910):

for the first part, he just plugged in t = 0 and evaluated H(t) = 25*cos( (pi/14)*t ) + 31 H(0) = 25*cos( (pi/14)*0 ) + 31 H(0) = 25*cos( 0 ) + 31 H(0) = 25*1 + 31 H(0) = 25 + 31 H(0) = 56

OpenStudy (anonymous):

What about the second part? Specifically, (pi/14)t = 2pi => t = 28 seconds

jimthompson5910 (jim_thompson5910):

well I would do it a different way (but you get the same answer) In general, y = A*cos(Bt - C) + D has a period of T = 2pi/B So... T = 2pi/B T = 2pi/(pi/14) T = 2pi*(14/pi) T = 28pi/pi T = 28 So the period of this function is 28 seconds. It will take 28 seconds to do a complete revolution.

OpenStudy (anonymous):

Ok this was very helpful. Thank you.

jimthompson5910 (jim_thompson5910):

you're welcome